OpenStudy (anonymous):

Factor the expression completely. 49x^2-81 How do i do thiss?!

jimthompson5910 (jim_thompson5910):

Hint: Use the difference of squares formula \[\Large a^2 - b^2 = (a-b)(a+b)\]

OpenStudy (anonymous):

uhmmm.. what? hahaha

OpenStudy (raden):

try this : a^2b^2 - c^2 = (ab + c)(ab - c)

jimthompson5910 (jim_thompson5910):

Similar example \[\Large 36x^2 - 25\] \[\Large (6x)^2 - 5^2\] \[\Large (6x-5)(6x+5)\] So \[\Large 36x^2 - 25\] factors to \[\Large (6x-5)(6x+5)\]

OpenStudy (anonymous):

so for mine, it would be 7x^2-9^2?

jimthompson5910 (jim_thompson5910):

close, more like (7x)^2 - 9^2

jimthompson5910 (jim_thompson5910):

keep going to get ...???

OpenStudy (anonymous):

(7x-9)(7x+9)

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

idk what to do for the next part

jimthompson5910 (jim_thompson5910):

you're done, it's factored completely

OpenStudy (anonymous):

so that's the answer?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

(7x-9)(7x+9)?? Okay thanks!

jimthompson5910 (jim_thompson5910):

yes, you're welcome